OpenStudy (lilai3):

Please help! Will award medal! Ms. Parker charges $29.95 for a bouquet of one dozen roses. Last year, she paid her supplier $4.50- per dozen roses. This year, she paid $3.25 more per dozen. How much less profit did she make this year on 20 dozen bouquets?

OpenStudy (anonymous):

One step at a time: How much of a profit did she make last year on a dozen roses?

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

total profit last year = 29.95-4.5 total profit this year = 29.95-(4.5+3.25) by subtracting you get the answer for a dozen multiplying by 20 you get the answer

OpenStudy (lilai3):

wait sorry it took so look, qwe,qwe uhm it was the connection.. wait i still don't get it. the question is asking, 'how much less profit did she make this year on 20 dozen bouqets?'

OpenStudy (anonymous):

Right. I read the same thing. So you have to figure out how much profit she made last year on 1 dozen. Then you have to figure out how much profit she made this year on 1 dozen. The difference between them is how much less she made this year than last year on 1 dozen. Multiply that by 20 and you have your answer.

OpenStudy (anonymous):

its $65 or 12.77% less profit

OpenStudy (lilai3):

oh so it's simple. you just do 3.25 times 20 and you get 65

OpenStudy (anonymous):

Yes!! :-)

OpenStudy (anonymous):

absolutely

OpenStudy (lilai3):

oh i get it now. thank you BOTh. i really wish i could award two medals, but since i cannot, i will fan one of you and give a medal to the other at random

OpenStudy (anonymous):

No problem!

OpenStudy (lilai3):

would you be willing to help me on two more of theses problems?